OpenStudy (anonymous):

Solve this system of equations using substitution. Put your answer in ordered pair form, (x, y). 5x + y = 1 3x + 2y = 2

OpenStudy (anonymous):

y=1-5x 3x+2(1-5x)=2 3x+2-10x=2 3x-10x=2-2 -7x=0 x=0 5x+y=1 0+y=1 y=1 (0,1)